Can you help me with a substitution problem? I am still getting confused,usually with the second part of the problem.It's not in a book.I do not think there is a solution for both of them to make them come out correctly.I really need simple step by step help in order to learn how to do them on my own.Can I just plug each of the sets of numbers given? Your help is greatly appreciated. .The problem is:

x=2y-8 3x-2y= -12
The choices were: a)2,-3 b)0,4 c)-4,4 d)-2,3 e)none
I CHOSE E - none of them (I almost chose C)

Answer by checkley71(8403)

X=2Y-8
3X-2Y=-12
3(2Y-8)-2Y=-12
6Y-24-2Y=-12
4Y=-12+24
4Y=12
Y=12/4
Y=3 ANSWER.
X=2*3-8
X=6-8
X=-2 ANSWER.
PROOF:
3*-2-2*3=-12
-6-6=-12
-12=-12